# Real-Time Wayfinding Assistant for Blind and Low-Vision Users

### v1

@@ 4.4. Comparative Analysis and Key Observations
  • Accuracy vs. Efficiency Trade-off: While VLM and LLM-based approaches achieve competitive accuracy in some cases, their response times are significantly higher, making them less practical for real-time navigation. The proposed DFS-based method (Approach 6) provides a favorable trade-off, offering a substantial improvement in response time (approximately 3-10 times faster) while maintaining competitive accuracy (around 2-3 times better), without a severe loss in performance. This demonstrates the method’s ability to balance precision with efficiency, making it more suitable for real-time applications.
  • Environment-Specific Performance: The proposed method exhibits superior performance in outdoor scenes, achieving the lowest MAE and highest accuracy among all approaches. In particular, Approach 6 achieves a 106-152% improvement in accuracy and a 16-24% reduction in MAE compared to other methods, suggesting that the algorithm is particularly well-suited for more complex and variable environments like outdoor settings.
  • Model-Specific Observations: Among LLM-based methods, gemini-1.5-pro consistently outperforms other models (Phi-3.5-mini-instruct and Phi-3-mini-4k-instruct) in both accuracy and MAE. However, the increased computational cost associated with these models limits their real-time applicability. In contrast, the proposed DFS-based method achieves a similar or better performance in terms of accuracy, with a significantly lower response time (roughly 3-10 times faster), making it more suitable for practical deployment.
  • Scalability and Real-Time Suitability: The rapid processing times of the proposed method indicate its potential for deployment in real-time assistive navigation systems. Specifically, Approach 6 offers a 3-10 times faster response time than other methods, which is critical for time-sensitive applications. Furthermore, its robustness across both indoor and outdoor scenes suggests scalability to diverse environments, ensuring that the method remains effective in varying real-world conditions.
The DFS-based optimal pathfinding approach offers several key advantages that make it especially suitable for real-time navigation assistance for blind or visually impaired (BLV) individuals. Its lightweight DFS-based pathfinding algorithm demands significantly less computational power than conventional deep learning models, allowing efficient operation on mobile devices without high-end hardware. By utilizing depth variations, the method accurately distinguishes walkable paths from obstacles, enhancing path detection reliability. Additionally, it ensures real-time responsiveness by providing rapid path updates, which supports seamless interaction in dynamic environments. Furthermore, the system is both user-friendly and cost-effective, requiring only a single-camera smartphone and eliminating the need for additional sensors or specialized equipment.
@@ 5. Usability Study
To assess the usability of PathFinder and the Topline System, we conducted a study with 15 blind or visually impaired (BLV) participants from diverse age groups, education levels, mobility aids, and geographical locations. The study sought to answer critical questions such as:
@@
  • What were the biggest issues that participants encountered when using the app?
@@
Participants were chosen based on various blindness characteristics, such as congenital blindness, visual loss caused by medical disorders (diabetic retinopathy, glaucoma, cataract, etc.), and accidents. The sample also varied in age, education level, mobility aids, and technology knowledge, allowing for a more thorough review of the system. Table 7 provides detailed participant information, including their blindness profile, approach procedures, ease of use ratings, response times, and method rankings. Additionally, participants provided feedback on indoor (7 participants) and outdoor (8 participants) navigation experiences, using a five-point Likert scale, with results visualized through boxplots showing percentile distributions.
@@ 5.1. Participant Demographics and Mobility Aids
The study included 15 participants, aged 25 to 80 years, with varying causes of blindness, including birth conditions, accidents, diabetic retinopathy, glaucoma, strokes, cataracts, and macular degeneration. Educational backgrounds ranged from no formal education to university graduates, and participants were from both urban and rural areas across Bangladesh.
@@
Table 4 summarizes the participants’ information, including their age, source of blindness, education level, mobility aid utilized, and location.
@@ 5.2. Participant Engagement and Feedback Collection
To ensure a diverse evaluation context, we implemented various engagement strategies during the usability study of our application. Some engaged in conversations with hospital patients, shopkeepers, neighbors, and community members, while others requested feedback from teachers, friends, and family members. This approach provided insights into real-world navigation challenges and user preferences across different social settings. All participants were informed of the purpose of the testing and provided their consent to participate, ensuring that they understood their role in the evaluation process and the significance of their feedback.
@@ 5.3. Ease of Use and Response Time
Figure 5. Distribution of Participants’ Ease of Use Ratings.
Figure 6. Distribution of Participants’ Response Time for Providing Feedback on the App.
@@ 5.4. Participant Feedback on Indoor and Outdoor Navigation
Participants evaluated their navigation experience in indoor (7 participants) and outdoor (8 participants) settings using a five-point Likert scale, where 1 = Very Difficult and 5 = Very Easy. The responses were visualized through boxplots in figure 10, illustrating percentile distributions.
@@ 5.4.2. Outdoor Navigation
Outdoor navigation received higher Likert ratings, as participants benefited from open spaces, familiar routes, and natural audio cues. However, a few participants rated outdoor navigation due to traffic noise, moving obstacles, and surface irregularities.
Figure 7. Usability ratings (1–5) for navigation schemes across indoor (purple) and outdoor (grey) scenes. Narrow boxplots show scene-wise rating distributions per scheme.
@@ 5.6.1. User Insights on Methods
Participants provided qualitative feedback on why they appreciated or disliked specific strategies. Why Participants Preferred Method 6 (DFS-Based):
  • ”This way seems the most natural and accurate. I don’t have to stop and think; I simply move.”
  • ”It gives me clear paths quickly without making me confused.”
  • ”Works well in crowded places where I need fast, precise decisions.”
Figure 9. Bar chart showing the number of participants who selected each of the six approaches as their top preference for both indoor and outdoor scenes.
@@ 5.8. Learning the App & Adaptability
  • Table 6 demonstrates that 66.67% of participants learnt the app in less than a minute, demonstrating a straightforward design.
  • Users Who Did Not Need Training: These participants, particularly those who were familiar with assistive technology, considered the app easy to use.
  • Users Who Needed Training: Some users (particularly elderly ones) preferred hands-on instruction before feeling comfortable using the program.
@@ 5.9.1. What Worked Well
  • Diverse navigation needs: Mobility aids and environmental factors significantly influenced participant preferences.
  • Indoor navigation challenges: Indoor environments presented more difficulties, reflected in lower Likert ratings and higher response times.
  • Outdoor navigation advantages: Participants found outdoor navigation easier, though some faced challenges with environmental distractions.
  • Method 6 (DFS-Based) was the clear favorite: 80% of participants (12 out of 15 from table 7) selecting it as their top priority due to various advantages such as good accuracy, speed, and ease of use.
@@ 5.9.2. Challenges Identified
  • Some users had difficulty interpreting methods: A few participants found certain navigation strategies, particularly In Degree (DOF 180), to be very complex.
@@
  • Response time variability: While the app functioned well in most cases, some users reported slight delays in direction updates when obstacles changed rapidly.
  • Request for haptic feedback: Users suggested incorporating haptic feedback, as voice output may not be practical in noisy or crowded places.
@@
Table 7. Demographic details, mobility aids, approach procedures, ease of use, response times, and method rankings from participant feedback in the usability study.
@@
13 Tunnel Vision Outdoor 50 secs 8 minutes In Degree ¿ Clock Hour Hand ¿ Left/Forward/Right I liked method 6 because it felt natural. The degree-based method was interesting but sometimes required stopping to process.
@@
6. Limitations and Potential Future Work
While our suggested navigation system provides an efficient and cost-effective solution for blind and low-vision (BLV) users, some limitations must be addressed in future study.
Another limitation is working in low-light circumstances, as the system is based on image-based depth estimate. Depth accuracy diminishes in dark situations, making obstacle detection less accurate. Future enhancements could include the use of infrared or LiDAR-based sensors to increase performance in low-light conditions.
Our technique doesn’t struggle as much in dynamic situations, where moving objects like pedestrians and automobiles present unforeseen challenges. Currently, the technology is intended for static scenarios, which limits its real-time adaptability. However, using haptic feedback instead of voice feedback might help resolve this issue, enabling more dynamic and responsive interaction. Future editions could investigate the use of real-time object tracking and motion prediction algorithms to deal with dynamic impediments more efficiently.
By addressing these limitations, we hope to improve our system’s stability, usability, and real-world application, resulting in more independent mobility for BLV individuals.
7. Conclusion
This study presented a novel, computationally efficient navigation system designed to enhance mobility for Blind and Low Vision (BLV) individuals. By leveraging a depth-based pathfinding algorithm with a Depth-First Search (DFS) approach, we demonstrated a low-cost, real-time solution that accurately detects free paths in both indoor and outdoor environments. Comparative evaluations showed that our proposed system offers a favorable trade-off between accuracy and computational efficiency, making it a viable alternative to existing AI-powered navigation solutions that require high-end hardware.
Our usability study, conducted with BLV participants, highlighted the system’s effectiveness in real-world settings, with the majority of users preferring our approach for its intuitive directional guidance and low inference time. While outdoor navigation yielded higher satisfaction due to fewer obstacles and broader spaces, indoor navigation posed challenges that require further refinement.

### v2

@@ 4.5. Comparative Analysis and Key Observations
Accuracy vs. Efficiency Trade-off. While VLM and LLM-based approaches achieve competitive accuracy in some cases, their response times are significantly higher, making them less practical for real-time navigation. The proposed pathfinding method (Approach 6) provides a favorable trade-off, offering a substantial improvement in response time (approximately 3-10 times faster) while maintaining competitive accuracy (around 2-3 times better), without a severe loss in performance. This demonstrates the method’s ability to balance precision with efficiency, making it more suitable for real-time applications.
Environment-Specific Performance. The proposed method exhibits superior performance in outdoor scenes, achieving the lowest MAE and highest accuracy among all approaches. In particular, Approach 6 achieves a 106-152% improvement in accuracy and a 16-24% reduction in MAE compared to other methods, suggesting that the algorithm is particularly well-suited for more complex and variable environments like outdoor settings.
Model-Specific Observations. Among LLM-based methods, gemini-1.5-pro consistently outperforms other models (Phi-3.5-mini-instruct and Phi-3-mini-4k-instruct) in both accuracy and MAE. However, the increased computational cost associated with these models limits their real-time applicability. In contrast, the proposed pathfinding method achieves a similar or better performance in terms of accuracy, with a significantly lower response time (roughly 3-10 times faster), making it more suitable for practical deployment.
Scalability and Real-Time Suitability. The rapid processing times of the proposed method indicate its potential for deployment in real-time assistive navigation systems. Specifically, Approach 6 offers a 3-10 times faster response time than other methods, which is critical for time-sensitive applications. Furthermore, its robustness across both indoor and outdoor scenes suggests scalability to diverse environments, ensuring that the method remains effective in varying real-world conditions.
To sum up, the pathfinding approach offers several key advantages that make it especially suitable for real-time navigation assistance for blind or visually impaired (BLV) individuals. Its lightweight pathfinding algorithm demands significantly less computational power than conventional AI models, allowing efficient operation on mobile devices without high-end hardware. By utilizing depth variations, the method accurately distinguishes walkable paths from obstacles, enhancing path detection reliability. Additionally, it ensures real-time responsiveness by providing rapid path updates, which support seamless interaction in dynamic environments. Furthermore, the system is both user-friendly and cost-effective, requiring only a single-camera smartphone and eliminating the need for additional sensors or specialized equipment.
@@ 5. Usability Study
To assess the usability of PathFinder and the Topline System, we conducted a study with 15 blind or visually impaired (BLV) participants from diverse age groups, education levels, mobility aids, and geographical locations. The study sought to answer the following critical questions:
@@
  • What were the biggest issues that participants encountered while using the app?
@@ 5.1. Participant Demographics and Mobility Aids
The study included 15 participants, aged 25 to 80 years, with varying causes of blindness, including birth conditions, accidents, diabetic retinopathy, glaucoma, strokes, cataracts, and macular degeneration. Educational backgrounds ranged from no formal education to university graduates, and participants were from both urban and rural areas across Bangladesh (Table 4).
@@ 5.2. Participant Engagement and Feedback Collection
To ensure a diverse evaluation context, we implemented various engagement strategies during the usability study of our application. We engaged in conversations with hospital patients, shopkeepers, neighbors, and community members. This approach provided insights into real-world navigation challenges and user preferences across different social settings. All participants were informed of the purpose of the testing and provided their consent to participate, ensuring that they understood their role in the evaluation process and the significance of their feedback.
@@ 5.4. Participant Feedback on Indoor and Outdoor Navigation
Participants evaluated their navigation experience in indoor (7 participants) and outdoor (8 participants) settings using a five-point Likert scale, where 1 = Very Difficult and 5 = Very Easy. The responses were visualized through boxplots in figure 10.
@@ 5.4.2. Outdoor Navigation
Outdoor navigation received higher Likert ratings, as participants benefited from open spaces, familiar routes, and natural audio cues. However, a few participants rated outdoor navigation low due to traffic noise, moving obstacles, and surface irregularities.
Figure 8. Usability ratings (1–5) for navigation schemes across indoor (purple) and outdoor (grey) scenes.
@@ 5.6.1. User Insights on Methods
Participants provided qualitative feedback on why they appreciated or disliked specific strategies. Why Participants Preferred Method 6 (Pathfinder Algorithm):
  • This way seems the most natural and accurate. I don’t have to stop and think; I simply move.
  • It gives me clear paths quickly without making me confused.
  • Works well in crowded places where I need fast, precise decisions.
@@ 5.8. Learning the App & Adaptability
  • Table 6 (Appendix B) shows that 66.67% of participants learned the app in under a minute, indicating a straightforward design.
  • Users Who Did Not Need Training. These participants, particularly those who were familiar with assistive technology, considered the app easy to use.
  • Users Who Needed Training. Some users (particularly elderly ones) preferred hands-on instruction before feeling comfortable using the program.
@@ 5.9.1. What Worked Well
  • Diverse navigation needs. Mobility aids and environmental factors significantly influenced participant preferences.
  • Indoor navigation challenges. Indoor environments presented more difficulties, reflected in lower Likert ratings and higher response times.
  • Outdoor navigation advantages. Participants found outdoor navigation easier, though some faced challenges with environmental distractions.
  • Method 6 (Pathfinder Algorithm) was the clear favorite. 80% of participants (12 out of 15 from table 8) selecting it as their top priority due to various advantages such as good accuracy, speed, and ease of use.
@@ 5.9.2. Challenges Identified
  • Some users had difficulty interpreting methods. A few participants found certain navigation strategies, particularly in Degree (DOF 180), to be very complex.
@@
  • Response time variability. While the app functioned well in most cases, some users reported slight delays in direction updates when obstacles changed rapidly.
  • Request for haptic feedback. Users suggested incorporating haptic feedback, as voice output may not be practical in noisy or crowded places.
@@
6. Limitations
Another limitation is working in low-light circumstances, as the system is based on an image-based depth estimate. Depth accuracy diminishes in dark situations, making obstacle detection less accurate. Future enhancements could include the use of infrared or LiDAR-based sensors to increase performance in low-light conditions.
Although our technique doesn’t struggle as much in dynamic situations, where moving objects like pedestrians and automobiles present unforeseen challenges. Our technology is intended for static scenarios, which limits its real-time adaptability. However, using haptic feedback instead of voice feedback might help resolve this issue, enabling more dynamic and responsive interaction. Future editions could investigate the use of real-time object tracking and motion prediction algorithms to deal with dynamic impediments more efficiently.
By addressing these limitations, we aim to improve our system’s stability, usability, and real-world application, resulting in more independent mobility for BLV individuals.
7. Conclusion
This study presented a novel, computationally efficient navigation system designed to enhance mobility for Blind and Low Vision (BLV) individuals. By leveraging a depth-based pathfinding algorithm, we demonstrated a low-cost, real-time solution that accurately detects free paths in both indoor and outdoor environments. Comparative evaluations showed that our proposed system offers a favorable trade-off between accuracy and computational efficiency, making it a viable alternative to existing AI-powered navigation solutions that require high-end hardware.
Our usability study, conducted with 15 BLV participants, highlighted the system’s effectiveness in real-world settings, with the majority of users preferring our approach for its intuitive directional guidance and low inference time. While outdoor navigation yielded higher satisfaction due to fewer obstacles and broader spaces, indoor navigation posed challenges that require further refinement.
Appendix
A. Application Design Decisions
This section summarizes the design of the haptic feedback mechanism used in the application. A clock-based directional feedback system was adopted for intuitive navigation, where each clock position corresponds to a specific vibration pattern. Short pulses represent directions from 12:00 to 3:00, while long pulses represent directions from 9:00 to 11:30. The mapping between clock time, vibration sequence, and vibration type is shown in Table 7.
Table 7. Vibration Patterns and Types for Clock-Based Directional Feedback
Clock Time Vibration Pattern (ms) Vibration Type
12:$0$ [0, 200] Single Short Pulse
12:30 [0, 200] Single Short Pulse
1:$0$ [0, 200, 100, 200] Double Short Pulse
1:30 [0, 200, 100, 200] Double Short Pulse
2:$0$ [0, 200, 100, 200, 100, 200] Triple Short Pulse
2:30 [0, 200, 100, 200, 100, 200] Triple Short Pulse
3:$0$ [0, 200, 100, 200, 100, 200, 100, 200] Quad Short Pulse
9:$0$ [0, 600, 100, 600, 100, 600] Triple Long Pulse
9:30 [0, 600, 100, 600, 100, 600] Triple Long Pulse
10:$0$ [0, 600, 100, 600] Double Long Pulse
10:30 [0, 600, 100, 600] Double Long Pulse
11:$0$ [0, 600] Single Long Pulse
11:30 [0, 600] Single Long Pulse
B. User Study
A usability study was conducted with 15 participants with different blindness profiles, mobility aids, and levels of experience. The study evaluated ease of use, response times, and preferences for navigation schemes. Tables 8 and 9 summarize participant demographics, interaction approaches, and user feedback.
Table 8. Demographic details, mobility aids, approach procedures, ease of use, response times, and method rankings from participant feedback.
@@ B. User Study
13 Tunnel Vision Outdoor 50 secs 8 minutes In Degree ¿ Clock Hour Hand ¿ Left/Forward/Right I liked method 6 because it felt natural. The degree-based method was interesting, but sometimes required stopping to process.